Protect Your Freedom to Write Programs Join the League for Programming Freedom

(Version of January 26, 1991)

Ten years ago, programmers were allowed to write programs using all the techniques they knew, and providing whatever features they felt were useful. This is no longer the case. The new monopolies, software patents and interface copyrights, have taken away our freedom of expression and our ability to do a good job.

“Look and feel” lawsuits attempt to monopolize well-known command languages; some have succeeded. Copyrights on command languages enforce gratuitous incompatibility, close opportunities for competition, and stifle incremental improvements.

Software patents are even more dangerous; they make every design decision in the development of a program carry a risk of a lawsuit, with draconian pretrial seizure. It is difficult and expensive to find out whether the techniques you consider using are patented; it is impossible to find out whether they will be patented in the future.

The League for Programming Freedom is a grass-roots organization of professors, students, businessmen, programmers and users dedicated to bringing back the freedom to write programs. The League is not opposed to the legal system that Congress intended—copyright on individual programs. Our aim is to reverse the recent changes made by judges in response to special interests, often explicitly rejecting the public interest principles of the Constitution.

The League works to abolish the new monopolies by publishing articles, talking with public officials, boycotting egregious offenders, and in the future may intervene in court cases. On May 24, 1989, the League picketed Lotus headquarters on account of their lawsuits, and then again on August 2, 1990. These marches stimulated widespread media coverage for the issue. We welcome suggestions for other activities, as well as help in carrying them out.

Membership dues in the League are $42 per year for programmers, managers and professionals; $10.50 for students; $21 for others. Please give more if you can. The League’s funds will be used for filing briefs; for printing handouts, buttons and signs; whatever will persuade the courts, the legislators, and the people. You may not get anything personally for your dues—except for the freedom to write programs. The League is a non-profit corporation, but not considered a tax-exempt charity. However, for those self-employed in software, the dues can be a business expense.

The League needs both activist members and members who only pay their dues. We also greatly need additional corporate members; contact us for information.

If you have any questions, please write to the League, phone (617) 243-4091, or send Internet mail to league@prep.ai.mit.edu.

Jack Larsen, President Chris Hofstader, Secretary Steve Sisak, Treasurer

Jack Larsen can be contacted at (708) 698-1160; Fax (708) 698-6221.

To join, please send a check and the following information to:

League for Programming Freedom 1 Kendall Square #143 P.O. Box 9171 Cambridge, MA 02139

(Outside the US, please send a check in US dollars from a bank with connections to the US, to save us check cashing fees.)

Your name:

The address where we should send League mailings (a few each year):

The company you work for, and your position:

Your phone numbers (home, work or both):

Your email address, so we can contact you for demonstrations or for writing letters. (If you don’t want us to contact you for these things, please say so, but please give us your email address anyway.)

Is there anything about you which would enable your endorsement of the LPF to impress the public? For example, if you are or have been a professor or an executive, or have written software that has a good reputation, please tell us.

Would you like to help with LPF activities?

The corporate charter of the League for Programming Freedom states:

The purpose of the corporation is to engage in the following activities:

1. To determine the existence of, and warn the public about restrictions and monopolies on classes of computer programs where such monopolies prevent or restrict the right to develop certain types of computer programs.

2. To develop countermeasures and initiatives, in the public interest, effective to block or otherwise prevent or restrain such monopolistic activities including education, research, publications, public assembly, legislative testimony, and intervention in court proceedings involving public interest issues (as a friend of the court).

3. To engage in any business or other activity in service of and related to the foregoing paragraphs that lawfully may be carried on by a corporation organized under Chapter 180 of the Massachusetts General Laws.

The officers and directors of the League will be elected annually by the members.